amcgill:

Solve the given system of equations. 2y=-x+9 3x-6y=-15

swinn:

Solve by Substitution : Solve equation [2] for the variable x [2] 9x = 2y + 19 [2] x = 2y/9 + 19/9Plug this in for variable x in equation [1] [1] 3•(2y/9+19/9) - 6y = -15 [1] - 16y/3 = -64/3 [1] - 16y = -64 // Solve equation [1] for the variable y [1] 16y = 64 [1] y = 4 By now we know this much : x = 2y/9+19/9 y = 4 Use the y value to solve for x x = (2/9)(4)+19/9 = 3 Solution :